Math 544 (Section 001)  University of South Carolina
D. Meade  Fall 1996

Homework Problems for Chapter 1 Matrices and Systems of Linear Equations

Section Page Problems
1.1  8 # 1, 2, 5, 6, 8, 13, 16^{∗}, 25, 28, 31, 34, 38, 40, 43^{∗}, 48
1.2  19 # 1, 2, 4, 10, 12^{∗}, 19, 22, 34, 35^{∗}, 39, 44
1.3  35 # 14, 17, 20, 26^{∗}, 30, 31, 33, 36, 41^{∗}, 44, 47
1.4  43 # 2, 7, 10^{∗}, 12, 13, 15, 20, 21^{∗}, 24
1.6  62 # 5, 8, 11, 16, 21, 34, 40, 42, 45^{∗}, 49, 50^{∗}, 51, 54, 57
1.7  71 # 10, 11, 17, 20, 22, 26^{∗}, 33, 34, 35, 37^{∗}, 41, 43, 45, 46, 50
1.8  81 # 4, 5, 9–12, 16, 17, 19, 20, 23, 30, 31^{∗}, 42, 46^{∗}, 48
1.10 103 # 9, 12, 16, 19, 22, 23, 25, 28, 30, 35, 38^{∗}, 40, 44, 45^{∗}
1.11 112 # 3, 8, 12, 16, 18^{∗}, 20, 22, 23, 29^{∗}
Review 114 Supplemental Computational Exercises

# 2, 3, 4, 6, 8, 9, 10, 11, 12, 14–18

Review 115 Supplemental Conceptual Exercises

# 1, 2, 3, 5, 7, 8, 10, 11, 12, 13, 14

\underline{Notes}

• Problems marked with an asterisk (^{∗}) will be collected and graded. In general, problems will be due the day following the lecture in which the section was discussed.

• These problems have been selected as the minimum amount of work that you should be doing for this course. Some problems will be straightforward computations; some will require the construction of a logical argument using the terminology and concepts discussed in class; some problems will require you to expand your knowledge (with given information) and apply this to a new setting.

• If you need more practice, there are many other problems in the text that are similar to the ones I have assigned.

• There are two sections of review problems at the end of each chapter. None of these problems will be collected for grading. These problems should be a good test of your understanding of the main points in the chapter. I suggest solving these problems as one part of your preparation for an exam.

• Be sure you know how to do all of these problems. Do not hesitate to ask questions – in class, by e-mail, or during office hours.